OpenStudy (anonymous):

Sheg??? 18. A customer went to a garden shop and bought some potting soil for $12.50 and 6 shrubs. If the total bill was $71.00, how much did each shrub cost if they all cost the same? a) Choose and label a variable for this situation. b) Write an equation using the variable that would answer the question. c) Solve your equation, and check your answer. d) Use the solution to write a sentence answer to the question in the word problem

OpenStudy (lilg132):

Potting soil cost $12.50 The total amount including 6 shrubs and potting soil was $71.00 so you do $71.00 - $12.50 = $58.50 this is the total price of 6 shrubs now to get the price of each shrub you divide this amount by 6 58.50 / 6 = $9.75 each shrub cost

OpenStudy (lilg132):

a) Total bill = tb x = soil y = shrubs

OpenStudy (lilg132):

b) Tb = x + 6y

OpenStudy (lilg132):

c) 71.00 = 12.50 + 6y 71.00 - 12.50 = 6y 6y = 58.50 6y / 6 = 58.50 / 6 y = $9.75 each shrub

OpenStudy (lilg132):

dont forget to click good answer :)